The inflation associated with the oil price shocks in the 1970s after OPEC restricted the supply of oil is an example of

demand-pull inflation due to a supply shock.

demand-pull inflation due to a demand shock.

cost-push inflation due to a demand shock.

cost-push inflation due to a supply shock.

If initial equilibrium real Gross Domestic Product (GDP) is $400 billion, MPC = 0.9, and autonomous investment increases $40 billion, equilibrium real Gross Domestic Product (GDP) will be

$800 billion.

$360 billion.

$440 billion.

$600 billion.
Answer #1

1)

cost-push inflation due to a supply shock.

because of low supply of oil, it led to higher crude prices resulting in higher inflation.

2)

multiplier impact=1/(1-MPC)=1/(1-0.9)

=10

Incremental GDP due to higher investment=investment increases *multiplier impact=40*10=400 billion

equilibrium real Gross Domestic Product (GDP) will be=initial equilibrium real Gross Domestic Product (GDP)+Incremental GDP due to higher investment

=400+400

=800 billion

the above is answer..